OpenStudy (anonymous):

Draw a triangle with side lengths 3cm,4cm,and 5cm. There should be a right angle in your triangle .

jhonyy9 (jhonyy9):

yes because 3,4 and 5 are Pithagorean numbers what sign that 5^2 =3^2 +4^2 so 25=9+16 25=25